OpenStudy (anonymous):

y = 5x + 6 y = −x − 7 Which description best describes the solution to the system of equations?

pooja195 (pooja195):

Use the substitution method E1 tells you what y equals 5x+6=-x-7 Solve for x

OpenStudy (anonymous):

I got x = -13/6 but I'm still unsure of what the answers mean. A.Line y = 5x + 6 intersects line y = −x − 7. B.Lines y = 5x + 6 and y = −x − 7 intersect the x-axis. C.Lines y = 5x + 6 and y = −x − 7 intersect the y-axis. D.Line y = 5x + 6 intersects the origin.

pooja195 (pooja195):

@jim_thompson5910 can you help? Please? After seeing the answer choices i confused..

pooja195 (pooja195):

*im

jimthompson5910 (jim_thompson5910):

solve `5x+6=-x-7` for x to get 5x+6=-x-7 5x+6+x=-x-7+x 6x+6 = -7 6x+6-6= -7-6 6x = -13 6x/6 = -13/6 x = -13/6 so I agree with you @lilly249 I'm getting the same x value

jimthompson5910 (jim_thompson5910):

now plug x = -13/6 into either equation to find y y = 5x+6 y = 5(-13/6)+6 y = -65/6+6 y = -65/6+6*(6/6) y = -65/6+36/6 y = (-65+36)/6 y = -29/6

jimthompson5910 (jim_thompson5910):

the two lines cross at (-13/6, -29/6) this point is NOT on the x axis this point is NOT on the y axis this point is in quadrant Q3

jimthompson5910 (jim_thompson5910):

Choice B,C,D are wrong because of what I said in my last post

OpenStudy (anonymous):

How is it not in the x and y- axis? I'm confused, I'm sorry T_T

jimthompson5910 (jim_thompson5910):

if it were on the x axis, then the y coordinate would be 0 if it were on the y axis, then the x coordinate would be 0 if it were on the origin, then both x and y coordinates would be 0

OpenStudy (anonymous):

OHHHHHH ok thanks!!

jimthompson5910 (jim_thompson5910):

no problem